[10] 7. Use the Annihilator method to find a particular solution of the equation y – 4y + 4y = 2e24
Answer #1

Answer: - Given equation 27 y-uy+ 4y = 20 (0-2) y = 20 20 where o=d dac This is non- homogeneous linear differential equatio2 2 2x Co,+Cox) e 2x + t.czne y (x) + Up (x) { using @] me 24 Y(x) Yp (oc) = xe so particular solution of eqno is 4p(x)=xeza